OpenStudy (anonymous):

Prove the Identity. cos(4x)=cos^(4)x-6sin^(2)xcos^(2)x+sin^(4)x

OpenStudy (anonymous):

\[\cos^{4}x-6\sin^{2}xcos^{2}x+\sin^{4}x\] \[= .5(1+\cos^{2}(2x)) -6*0.5(1-\cos{2x})*0.5(1+\cos{2x}) +0.5(1-\cos{2x}) \] \[=.5+(.5\cos^{2}{2x})-3*0.5+(3*0.5\cos{2x})-3*0.5-(3*0.5\cos{2x})\] \[=-6.5+0.5\cos^{2}{2x}\] \[=-6.5+\frac{1}{4}(1+\cos4x)\] This is all i know .. it maybe further simplified . regards

OpenStudy (anonymous):

I have replaced 1 with: \[\large \sin^2(4x) + \cos^2(4x) = 1\]

OpenStudy (anonymous):

And somehow I am getting -8 and not -6 in the middle @rosiepose99 Check your question once again..

OpenStudy (anonymous):

I think I have done something wrong..

OpenStudy (anonymous):

it is -6

OpenStudy (anonymous):

\(\cos(4x) = \cos2(2x) \implies 1 - 2\sin^2(2x) \implies 1 - 2(\sin(2x)^2)\) \(= 1 - 2(2\sin(x)\cos(x))^2 \implies 1 - 2(4\sin^2(x) \cos^2(x)) \implies1 - 8\sin^2(x) \cdot(\cos^2(x))\) Now using: \[\cos^4(x) + \sin^4(x) = 1 - 2\sin^2(x) \cdot \cos^2(x)\] \[\implies 1 = \cos^4(x) + \sin^4(x) + 2\sin^2(x) \cdot \cos^2(x)\] Replace 1 with it: \[\implies1 - 8\sin^2(x) \cdot(\cos^2(x))\] \[\implies \cos^4(x) + \sin^4(x) + 2\sin^2(x) \cdot \cos^2(x) - 8\sin^2(x) \cdot \cos^2(x)\] \[ \implies \color{green}{\cos^4(x) + \sin^4(x) - 6\sin^2(x) \cdot \cos^2(x)}\]

OpenStudy (anonymous):

where you got cos^4x +sin^4x =1-2sin^2c*cos^2x why?

OpenStudy (anonymous):

Do you know : \[\cos^2(x) + \sin^2(x) = 1\] @rosiepose99

OpenStudy (anonymous):

And do you know: \[(a + b)^2 = a^2 + b^2 + 2ab\]

OpenStudy (anonymous):

here put \(a = cos^2{x}\) and b = \(sin^2(x)\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

We will get: \[(\cos^2(x) + \sin^2(x))^2 = (\cos^2(x))^2 + (\sin^2(x))^2 + 2\sin^2(x)\cos^2(x)\] \[(\cos^2(x) + \sin^2(x))^2 = \cos^4(x) + \sin^4(x) + 2\sin^2(x)\cos^2(x)\] As: \(\cos^2(x) + \sin^2(x) = 1\) put this on LHS: \[\implies 1 = \cos^4(x) + \sin^4(x) + 2\sin^2(x)\cos^2(x)\] Getting @rosiepose99

OpenStudy (anonymous):

yes, but where you replace sin^2x +cos^2x=1 you would be left with another sin^2x+cos^2x right...and it shows sin^4x +cos^4x??????

OpenStudy (anonymous):

I replace on left hand side..

OpenStudy (anonymous):

See: \[(\cos^2(x) + \sin^2(x))^2 = \cos^4(x) + \sin^4(x) + 2\sin^2(x)\cos^2(x)\] \[(1)^2 = \cos^4(x) + \sin^4(x) + 2\sin^2(x)\cos^2(x)\]

OpenStudy (anonymous):

Getting now @rosiepose99 put this value in place of 1 there..

OpenStudy (anonymous):

I will show you where I have replaced.. Just look for colored one: \[(\color{green}{\cos^2(x) + \sin^2(x)})^2 = \cos^4(x) + \sin^4(x) + 2\sin^2(x)\cos^2(x)\]

OpenStudy (anonymous):

Do reply @rosiepose99

OpenStudy (anonymous):

ok but it looks like u r using cos^2x +sin^2x twice

OpenStudy (anonymous):

Where??

OpenStudy (anonymous):

one sec.

OpenStudy (anonymous):

See there is difference between: \[\large (\cos^2(x) + \sin^2(x))^2\] and : \[\large (\cos^2(x))^2 + (\sin^2(x))^2\]

OpenStudy (anonymous):

First one is: \[\large (1)^2 \implies 1\]

OpenStudy (anonymous):

And the second one is: \[\large \cos^4(x) + \sin^4(x)\]

OpenStudy (anonymous):

cos^4(x)+sin^4(x)=1−2sin^2(x)⋅cos^2(x) this is where i am confused.....?y?

OpenStudy (anonymous):

See I have used the above one in my way. So can you forget this?/ Just apply the identity that I have just given that: \[(1)^2 = \cos^4(x) + \sin^4(x) + 2\sin^2(x)\cos^2(x)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

See I don't want you to cram it.. I will surely show you how I did..

OpenStudy (anonymous):

Just reply fast @rosiepose99 Okay??

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

See: \[(a+b)^2 = a^2 + b^2 + 2ab\] Right??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Subtract both the sides by 2ab: \[(a+b)^2 - 2ab = a^2 + b^2 + 2ab - 2ab\] Right??

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Didn't get??

OpenStudy (anonymous):

So 2ab and -2ab on right hand side will get cancel: \[(a+b)^2 - 2ab = a^2 + b^2 + \cancel{2ab} - \cancel{2ab}\] Rght??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So you are left with: \[a^2 + b^2 = (a+b)^2 - 2ab\] right?? I have just exchanged RHS and LHS..

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Put a = \(cos^2(x)\) and b = \(sin^2(x)\) you will get: \[(\cos^2(x))^2 + (\sin^2(x))^2 = (\sin^2(x) + \cos^2(x)) - 2\sin^2(x) \cdot \cos^2(x)\] Right ?? Look carefully..

OpenStudy (anonymous):

Sorry I did wrong..

OpenStudy (anonymous):

y did square twice

OpenStudy (anonymous):

Wait..

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

k

OpenStudy (anonymous):

\[(\cos^2(x))^2 + ((\sin^2(x))^2 = (\sin^2(x) + \cos^2(x))^2 - 2\sin^2(x) \cdot \cos^2(x)\] Just same as: \[a^2 + b^2 = (a+b)^2 - 2ab\]

OpenStudy (anonymous):

So firstly solve LHS: \[(\cos^2(x))^2 + (\sin^2(x))^2 = \cos^4(x) + \sin^4(x)\] Right?/ this is LHS..

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can u show/explain to me how/why you got cos4x=1-8sin^2xcos^2x

OpenStudy (anonymous):

Now solve RHS: \[(\sin^2(x) + \cos^2(x))^2 - 2\sin^2(x) \cdot \cos^2(x) \implies (1)^2 - 2\sin^2(x) \cdot \cos^2(x)\]

OpenStudy (anonymous):

on the LHS

OpenStudy (anonymous):

Do you know: \[\cos(2x) = 1-2\sin^2(x)\]

OpenStudy (anonymous):

no, is that the double angle identity?

OpenStudy (anonymous):

yeah it is, ok

OpenStudy (anonymous):

They are half angle and double identities.. You have to learn them..

OpenStudy (anonymous):

yes i remember

OpenStudy (anonymous):

Double angle..

OpenStudy (anonymous):

In LHS whatever be the angle on RHS it becomes half.. Okay??

OpenStudy (anonymous):

k

OpenStudy (anonymous):

So can I write: \(cos(4x)\) as \(cos2(2x)\) tell me??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So: \[\cos(4x) = 1 - 2\sin^2(2x)\] Right??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but then you squared,,,y

OpenStudy (anonymous):

Where??

OpenStudy (anonymous):

I am showing you step by step so be cool.. Last step you got??

OpenStudy (anonymous):

\[\cos(4x) = 1 - 2\sin^2(2x)\] Go till here??

OpenStudy (anonymous):

ok, i misread because you rewrote the power to the end of the parenthesis.

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

See: \(sin^2(x)\) can be written as : \((sin(x))^2\) Okay??

OpenStudy (anonymous):

Do you know this: \[\sin(2x) = 2 \sin(x)\cos(x)\]

OpenStudy (anonymous):

no i didn't,

OpenStudy (anonymous):

Remember this always..

OpenStudy (anonymous):

It is also double and half angle identity...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Do you want its proof or you will remember this??

OpenStudy (anonymous):

I just looked at notes and see the identity, thanks anyway

OpenStudy (anonymous):

\[1 - 2(\sin(2x)^2) \implies 1 - 2(2\sin(x)\cos(x))^2\] Right??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Square the last: \[1 - 2(2\sin(x)\cos(x))^2 \implies 1 - 2(4\sin^2(x)\cos^2(x))\] Right??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now multiply 2 with 4: \[1 - 2(4\sin^2(x)\cos^2(x)) \implies 1 - 8\sin^2(x)\cos^2(x)\]

OpenStudy (anonymous):

Got all the steps?? Form now replace 1 with that I told you above.. Got??

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Go through all the steps you will definitely get that..

OpenStudy (anonymous):

ok, back to the one you explained first we r replacing 1, but where did you get the cos^4x +sin^4x. Is it from the other side?

OpenStudy (anonymous):

why did you choose to cos^4x + sin^4x